OpenStudy (anonymous):

How do you factor x^(3)+125 ?

OpenStudy (anonymous):

Use wolframalpha.com to do it for you. But also (x^3 + y^3) = (x + y)(x^2 - xy + y^2). Set y to be 5 to get your answer

OpenStudy (anonymous):

loser

OpenStudy (anonymous):

i need to know how to do it, it ust shows the answer.

OpenStudy (anonymous):

Did you figure out how to do it? (x^3 + y^3) = (x + y)(x^2 - xy + y^2), and so plug 5 in for y to get (x^3 + 5^3) = (x + 5)(x^2 - x*5 + 5^2)

OpenStudy (anonymous):

no i didnt cause that website just gives you the answer. but thanks